# Sparse Attention Vectors: Generative Multimodal Model Features Are Discriminative Vision-Language Classifiers

### v2

@@ 4 Evaluation
We apply SAVs to two state-of-the-art LMMs—LLaVA-OneVision [41] and Qwen2-VL [91]. We also do a rigorous comparison of our method to strong few-shot and finetuning baselines on a variety of different discriminative vision-language tasks covering safety, VQA, and classification.
@@ 4.1 Implementation Details
We implemented our approach in PyTorch [70]. We use the official implementations of each model, and all of our experiments can be run on a single NVIDIA A100 GPU. More details of the implementation is included in the supplementary material in Section B.
4.2 Models
In our work we demonstrated the effectiveness of Sparse Attention Vectors utilizing the following models: (1) Qwen2-VL [90] improves on its predecessor with dynamic processing of images of varying resolutions into different numbers of visual tokens in order to align with human perception of those images. This dynamic processing paired with novel positional embeddings that effectively fuse positional information across modalities affords Qwen2-VL its SOTA performance on a variety of image-text and video-text tasks. (2) LLaVA-OneVision [42] is an open source LMM that performs well in single-image, multi-image and video tasks thanks to its pretraining and AnyRes visual processing.
Both Qwen2-VL and LLaVA-OneVision finetune on a variety of tasks, from visual question-answering to document understanding and video tasks. These much larger and diverse finetuning regimes contribute significantly to these models’ remarkable performance on an eclectic variety of generative vision-language tasks.
4.3 Datasets
@@
Image Classification. (1) EuroSAT [18] is a dataset that tackles the challenge of land use and land cover classification from satellite images. (2) Oxford-IIIT-Pets [69] is a dataset containing 37 categories of pet and roughly 200 images for each category. It is a very challenging dataset as it contains some breeds of cats that is very hard to discern from one another.
@@ 4.4 Baselines
For our primary results, we utilized SAVs with 20 examples per label. We compared our method with multiple state-of-the-art baselines. Zero-shot (ZS) baselines are implemented through querying the model directly and generating a response. In addition to ZS, we also compare to several test-time and finetuning few-shot methods (all with the exact same sample complexity as SAVs). For instance, we compare to the current state-of-the-art multimodal few-shot method, MTV [26] as well as LoRA [25] finetuning the model for each task. Since our method involves sparse, informed feature selection, we also compare to common naive feature extraction methods: using last-layer and all-layer attention vectors instead of selecting task’s specific heads.
@@
4.5 Results
Results are shown in Table 1. An advantage of our method is its adaptability to any discriminative vision-language task that has image, text, or interleaved image-text inputs. As such, we demonstrate that applying SAVs to a wide range of tasks in safety, VQA, and image classification outperforms all zero-shot baselines, and even drastically closing the gap with discriminative VLMs on image classification compared to SigLIP and CLIP. It is worth noting that these models cannot be directly compared to on tasks requiring interleaved image-text data. Furthermore, our approach even significantly improves over both state-of-the-art few-shot and finetuning methods. Finally, not only is our method broadly successful across a wide-range of tasks, but it also improves on challenging perception tasks that require visual and compositional reasoning abilities (e.g. BLINK and NaturalBench) that all VL models struggle with. For more results, please refer to our Supp. section in Section A.
In the following subsections, we demonstrate important properties and capabilities of our method through various ablation studies and additional experiments.
@@ 4.6 Ablations
We perform a comprehensive ablation study of our method on MHaluBench, NaturalBench, and EuroSAT (see Table 2). For more ablations, please refer to Section A.1 in the Supp. For all ablations, we use LLaVA-OneVision-7B.
Varying number of examples. In  Figure 4 (left), we examine the impact of varying the number of few-shot examples used in our method. Our primary results in Table 1 indicate that just 20 examples per label are necessary to yield state-of-the-art performance on a variety of discriminative VL tasks. This ablation shows that accuracy on these tasks can scale with increasing numbers of examples per label.
Varying number of attention vectors. Our method involves selecting a very sparse set of heads out of hundreds from which to extract attention vectors. We vary the number of attention vectors selected in Step 2 of our method and demonstrate that just 20 vectors are enough to realize nearly all of the classification accuracy of our method. Results are shown in Figure 4 (right).
@@
Comparing heads vs. layers. Based on prior work and transformer-architecture intuition, we treat the attention vectors outputted by the heads as a viable set of features for discriminative VL tasks. We verify this intuition by comparing the performance of selecting 2 sparse layers to selecting sparse heads as feature maps for our tasks. As shown in Table 2(b), head-based attention vectors outperform concatenated layer features on all three benchmarks.
Token position selection. Because the last-token of a sequence in a decoder-only LMM attends to all of the prior tokens in an input sequence, it is natural to extract SAVs from the heads of the last token. However, to validate this intuition, we compare the performance SAVs to extract sparse vectors from other tokens (first, middle, and last). Overall, our results in Table 2(c) show that the last token is the best option for selecting heads for SAVs.
@@ 4.7 Additional Experiments
(c) Impact of Token Position
MHB NB ES
Last 80.8 35.1 86.7
Middle 49.8 2.4 82.7
First 49.4 0 24.9
@@
Table 2: SAVs Ablations. We perform several ablations to identify the important aspects of our method that contribute to its effectiveness. In particular, we compare the effectiveness of (a) different classification methods, (b) head feature sparsification versus layer feaure sparsification, and (c) token positions from which to extract SAVs. Note: MHB represents MHaluBench, NB represents NaturalBench Group Score, and ES represents EuroSAT. For more ablations, please refer to Section A.1 in the Supplementary.
@@
In this subsection, we present experiments that demonstrate additional properties and capabilities of SAVs, beyond its use as features for discriminative VL tasks. Additional experiments can be found in Section A.2 of the Supplementary. For all experiments, we use LLaVA-OneVision-7B.
@@
Visualizing SAVs. SAVs are both an efficient and interpretable method for leveraging generative LMMs for discriminative VL tasks. To emphasize this point, we demonstrate the selected heads for hallucination detection, relative depth, and image classification in the first row of Figure 3. The visualizations demonstrate both the sparsity and specificity of the SAVs that are used for each individual task. Unlike other black-box methods, our approach clearly outlines exactly where in the model’s activation space informative attention vector features are extracted from. We furthermore show that the features extracted from these heads are useful for the given task in the second row of the figure, where we visualize the features outputted by the top selected head via t-SNE [22]. The fairly clear clustering of examples of the same label indicates that even with a single head, high-quality features are being selected as SAVs.
@@
Evaluating SAVs generalizability. Here, we ask whether the sparse heads SAVs heads extracted from one task, can generalize to another similar task. We utilize SAV heads from MHaluBench to evaluate on VLGuard and vice versa. We do a similar approach with LoRA, by swapping LoRA finetuned weights when evaluating on MHaluBench and VLGuard. Interestingly, our results in Table 3(a) show that heads extracted for SAVs generalize between tasks, while LoRA weights, as expected, overfit to the finetuned task.
@@
Comparing SAVs to CLIP/SigLIP on interleaved image-text tasks. As discussed in Section 1, SAVs are fully multimodal features able to represent inputs that are image-only, text-only, and even interleaved image-text. This is something that is not possible to directly replicate with CLIP and SigLIP models which have separate image and text encoders. Nevertheless, we compare our method to both CLIP and SigLIP on tasks that require interleaved image-text inputs. While SAVs can do this natively, we enable this comparison by concatenating the separate image and text features of the discriminative VLMs in order to evaluate on MHaluBench and NaturalBench. We find that our method vastly outperforms concatenated CLIP and SigLIP features on both benchmarks as shown in Figure 3(b). This result demonstrates the adaptability of our method to any discriminative VL task regardless of the input’s modality.
(a) Generalization
MHB VLG
Zero-Shot 34.7 31.4
MHB LoRA 34.5 33.9
MHB SAV 67.0 92.3
(b) Interleaved Tasks
MHB NB
CLIP 51.9 1.2
SigLIP 48.6 1.2
SAVs 80.8 35.1
Table 3: Additional SAVs Experiments. We perform additional experiments that (a) demonstrate the generalization of SAV heads to similar tasks as well as (b) show the effectiveness of SAV for tasks with interleaved image-text inputs.
@@ 5 Conclusion
Our research demonstrates the effectiveness of extracting Sparse Attention Vectors (SAVs) from the heads of an LMM and utilizing them directly for discriminative classification. Our method stands out by using only few-shot examples per label and only less than 1% of the heads to outperform zero-shot, few-shot, and fine-tuned baselines on a variety of VL tasks. In addition, SAVs allows generative LMMs to close the gap with discriminative VLMs on image classification tasks while also being an interpretable method that can generalize to similar tasks. Our ablations reveal the flexibility of using any classification method as a sparsification method for attention vectors and also shows that features are found as outputs of heads rather than layers. Overall, these results demonstrate that SAVs is a lightweight, performant, and generalizable method for extending generative LMMs’ capabilities to discriminative tasks. We are encouraged by the outcomes, and anticipate many directions for future work. In addition to methodological improvements, we look forward to the application of SAVs as features for multimodal retrieval, data compression, or more generally as a distilled representation for downstream models.
@@ 6 Limitations
Sparse Attention Vectors are a significant step in generalizing the capabilities of generative LMMs to discriminative tasks. Nevertheless, it is valuable to consider certain limitations of our approach. SAVs are a method that requires access to the model’s internal architecture and so may not be directly applicable to closed-source models like GPT-4 [67] and Gemini [83, 82]. Additionally, some tasks like image-text retrieval [84, 24] can benefit from more fine-grained confidence metrics attached to each label than proportion of voting heads per label. These challenges prompt future work in these directions as well as exciting questions about how to use SAVs as feature embeddings for other tasks.
@@ Acknowledgements.
Here we provide additional information about additional experimental results, qualitative examples, implementation details, and datasets. Specifically, Section A provides more experiment results, Section B provides additional implementation details, and Section C provides qualitative visualizations to illustrate our approach.
Appendix A Additional Experiment Results
We begin by presenting several additional ablations (Section A.1) that further demonstrate the benefits of our SAVs approach. We also present additional results (Section A.2) on BLINK Splits.
A.1 Additional Ablations
@@
A.2 Additional Results
(a) SAVs vs Few-Shot SigLIP
NB ES
SigLIP Few-Shot 1.2 88.2
SAVs 35.1 86.7
@@ A.2 Additional Results
Table 5: SAV Additional Results. We perform several additional experiments to demonstrate different properties and capabilities of SAVs. In particular, we evaluate the effectiveness of our method (a) compared to few-shot SigLIP, (b) on language-only tasks, and (c) when using it in an online learning setting. Note: MHB represents MHaluBench, NB represents NaturalBench Group Score, ES represents EuroSAT, and O.L. represents online learning.
@@
Comparing SAVs to few-shot SigCLIP. As SAVs are extracted with few-shot examples, we seek to compare our method to ton an analogous version of few-shot SigLIP. However, because SigLIP cannot be directly made few-shot, we adapt SigLIP as a few-shot class centroid classifier to make a fair comparison. In particular, we aggregate discriminative VLM embeddings into a mean embedding for each label. Then, just as in SAVs, we perform class centroid classification for each query using our set of mean SigLIP embeddings. We note that for an image classification task like EuroSAT, only image embeddings are necessary, but for MHaluBench multimodal embeddings are necessary. SAVs are inherently multimodal and so can be flexibly applied to both, but discriminative VLMs only have image-only or text-only embeddings. To work around this limiatation, we use SigLIP image features for EuroSAT and concatenate image and text features for MHaluBench. Interestingly, while SiGLIP is comparable to SAVs on EuroSAT, our method vastly outperforms SigLIP in the few-shot setting for MHaluBench. This suggests the generalizability of our method for both visual and more inherently multimodal tasks that discriminative VLMs struggle with. Our results are shown in Table 5(a).
@@
Appendix B Additional Implementation Details

### v3

@@ 4 Evaluation
We apply SAVs to two state-of-the-art LMMs—LLaVA-OneVision [44] and Qwen2-VL [97]. We also do a rigorous comparison of our method to strong few-shot and finetuning baselines on a variety of different image-text and image-only vision-language classification tasks covering safety, VQA, image-text retrieval, and simple classification.
@@ 4.1 Implementation Details
We implemented our approach in PyTorch [75]. We use the official implementations of each model, and all of our experiments can be run on a single NVIDIA A100 GPU. We also use the hyperparameters specified by the original model, which are all tuned for multi-task VL training such as the ones we report in our paper. More details of the implementation is included in the supplementary in Section 8.
Figure 3: Head Visualization. On the top row, we show the top-20 attention head locations for a given task. All the heads are indexed by their (head, layer) position, and the selected heads are highlighted. On the bottom row, we visualize the attention vector of few-shot examples for the top head of the given class with t-SNE clustering [25]. Each point represents the embedding of an input sample.
4.2 Benchmarks
@@
[…]
@@ 4.3 Baselines
For our results, we utilized SAVs with 20 examples per label. We compared our method with multiple SOTA baselines, including GPT-4o [72]. As a closed-source model, GPT-4o is presented simply as a strong closed-source baseline. Furthermore, we present the results of classic LMMs, LLaVA-1.5 and InstructBLIP, which have been previously applied to image classification [109]. Zero-shot (ZS) baselines are implemented by querying the model directly and generating a response. For image-text retrieval, ZS uses the SOTA generative scoring method VQAScore [56]. In addition to ZS, we also compare to several test-time and finetuning few-shot methods (all with the exact same sample complexity as SAVs). For instance, we compare to the current SOTA multimodal few-shot method, MTV [29] as well as doing 4-shot ICL and LoRA [28] finetuning for each task.
4.4 Results
@@
Results are shown in Table 1. An advantage of our method is its adaptability to any VL classification task that has image, text, or interleaved image-text inputs. Thus, we apply SAVs to a wide range of tasks in safety, VQA, image-text retrieval and image classification. Furthermore, our approach even significantly improves over SOTA ZS, few-shot, and finetuning methods. One interesting observation is that few-shot ICL consistently deteriorates performance. We hypothesize that instructioning tuning of LMMs disrupts the few-shot prompting capabilities gained during pretraining in favor of structured instructions as shown in [85, 13]. On the other hand, SAVs overcome this by directly operating on the internal activations. Finally, not only is our method successful across a wide-range of tasks, but it also improves on challenging visual perception and compositional reasoning tasks (e.g., BLINK and NaturalBench) that all VL models struggle with. Please refer to our supplementary section in Section 7 for mor results.
@@ 4.5 Ablations
We perform a comprehensive ablation study of our method on MHaluBench, NaturalBench, and EuroSAT (see Table 2). For more ablations, please refer to Section 7.1 in the Supp. For all ablations, we use LLaVA-OneVision-7B.
Varying number of examples. In  Figure 4 (left), we examine the impact of varying the number of few-shot examples used in our method. Our primary results in Table 1 indicate that just 20 examples per label are necessary to yield state-of-the-art performance on a variety of VL classification tasks. This ablation shows that accuracy on these tasks can scale with increasing numbers of examples per label.
Varying number of attention vectors. In order to extract attention vectors, we select a very sparse set of heads from hundreds. We vary the number of attention vectors selected in Step 2 of our method and demonstrate that just 20 vectors are enough to realize nearly all of the classification accuracy of our method. Results are shown in Figure 4 (right).
@@
Comparing heads vs. layers. Based on prior work and transformer-architecture intuition, we treat the attention vectors outputted by the heads as a viable set of features for classification tasks. We verify this intuition by comparing the performance of selecting 2 sparse layers to selecting sparse heads as feature maps for our tasks. As shown in Table 3(b), head-based attention vectors outperform concatenated layer features on all three benchmarks.
@@ 4.6 Additional Experiments
(a) Generalization
MHB VLG
Zero-Shot 34.7 31.4
VLG LoRA 34.1 90.0
VLG SAV 55.3 94.3
(b) Interleaved Tasks
MHB NB
CLIP 51.9 1.2
SigLIP 48.6 1.2
SAVs 80.8 35.1
(c) SAVs vs Few-Shot SigLIP
NB ES
SigLIP Few-shot Mean 1.2 88.2
SigLIP Few-shot Cross-Modal 6.8 86.4
SAVs 35.1 86.7
Table 2: Additional SAVs Experiments. We (a) demonstrate the generalization of SAV heads to similar tasks, (b) show the effectiveness of SAV for tasks with interleaved image-text inputs, and (c) compare SAVs with few-shot formulations of CLIP and SigLIP.
@@
Table 3: SAVs Ablations. We perform several ablations to identify the important aspects of our method that contribute to its effectiveness. In particular, we compare the effectiveness of (a) different classification methods and (b) head feature sparsification versus layer feature sparsification. Note: MHB represents MHaluBench, NB represents NaturalBench Group Score, and ES represents EuroSAT. For more ablations, please refer to Section 7.1 in the Supplementary.
@@
In this subsection, we present experiments that demonstrate additional properties and capabilities of SAVs, beyond its use as features for VL classification tasks. Additional experiments can be found in Section 7.2 of the Supplementary. For all experiments, we use LLaVA-OneVision-7B.
@@
Visualizing SAVs. SAVs are both an efficient and interpretable method for leveraging generative LMMs for VL classification tasks. To emphasize this point, we show the selected heads for hallucination detection, relative depth, and image classification in the first row of Figure 3. The visualizations demonstrate both the sparsity and specificity of the SAVs that are used for each task. Unlike prompting and finetuning methods, our approach clearly outlines exactly where in the model’s activation space informative attention vector features are extracted from. We furthermore show that the features extracted from these heads are useful for the given task in the second row of the figure, where we visualize the features outputted by the top selected head for each few-shot sample via t-SNE [25]. The clear clustering of examples of the same label indicates that even with a single head, high-quality features are being selected as SAVs.
@@
Evaluating SAVs generalizability. Here, we ask whether SAVs extracted from one task, can generalize to another similar task. We utilize SAV heads from VLGuard to evaluate on MhaluBench and vice versa. We do a similar approach with LoRA, by applying LoRA finetuned on VLGuard to MHaluBench. Interestingly, our results in Table 2(a) show that SAVs generalize between tasks, while LoRA weights, as expected, overfit to the finetuned task.
@@
Comparing SAVs to CLIP/SigLIP on interleaved image-text tasks. As discussed in Section 1, SAVs are fully multimodal features able to represent inputs that are image-only, text-only, and even interleaved image-text. This is something that is not possible to directly replicate with CLIP and SigLIP models which have separate image and text encoders. Nevertheless, we compare our method to both CLIP and SigLIP on tasks that require interleaved image-text inputs. While SAVs can do this natively, we enable this comparison by concatenating the separate image and text features of CLIP and SigLIP in order to evaluate on MHaluBench and NaturalBench. We find that our method vastly outperforms concatenated CLIP and SigLIP features on both benchmarks as shown in Figure 2(b). This result demonstrates the adaptability of our method to any VL classification regardless of the input’s modality.
Comparing SAVs to few-shot SigCLIP. As SAVs are extracted with few-shot examples, we compare our method to an analogous version of few-shot SigLIP. However, because SigLIP cannot be directly made few-shot, we adapt SigLIP as a few-shot class centroid classifier. One method used is the current SOTA few-shot classification method for encoder models Cross-Modal Adaptation [55]. We apply a second method where we aggregate CLIP/SigLIP embeddings into a mean embedding for each label. Then, just as in SAVs, we perform class centroid classification for each query using our set of mean SigLIP embeddings. We note that for image classification like EuroSAT, only image embeddings are needed, but for MHaluBench, multimodal embeddings are necessary. SAVs are inherently multimodal and so can be flexibly applied to both, but CLIP/SigLIP only have image-only or text-only embeddings. To overcome this, we use SigLIP image features for EuroSAT and concatenate image and text features for MHaluBench. Interestingly, while SiGLIP is comparable to SAVs on EuroSAT, our method vastly outperforms SigLIP in the few-shot setting for MHaluBench, suggesting generalizability of our method for a variety of multimodal tasks that CLIP-like struggle with. Our results are shown in Table 2(c).
@@ 5 Conclusion
Our research demonstrates the effectiveness of extracting Sparse Attention Vectors (SAVs) from the heads of an LMM and utilizing them directly for vision-language classification. Our method stands out by using only few-shot examples per label and only less than 1% of the heads to outperform zero-shot, few-shot, and fine-tuned baselines on a variety of image-text and image-only tasks. In addition, SAVs allows generative LMMs to close the gap with closed-source GPT-4o while also being an interpretable method that can generalize to similar tasks. Our ablations reveal the flexibility of using any classification method as a sparsification method for attention vectors and also shows that features are found as outputs of heads rather than layers. Overall, these results show that SAVs is a lightweight, performant, and generalizable method for extending generative LMMs’ multimodal classification abilities. We are encouraged by the outcomes, and anticipate many directions for future work. In addition to methodological improvements, we look forward to the application of SAVs as features for multimodal retrieval, data compression, or more generally as a distilled representation for downstream models.
@@ 6 Limitations
Sparse Attention Vectors are a significant step in generalizing the capabilities of generative LMMs to classification tasks. Nevertheless, it is valuable to consider certain limitations of our approach. SAVs are a method that requires access to the model’s internal architecture and so may not be directly applicable to closed-source models like GPT-4 [72] and Gemini [89, 88]. Additionally, some tasks like image-text retrieval [90, 27] can benefit from more fine-grained confidence metrics attached to each label than proportion of voting heads per label. These challenges prompt future work in these directions as well as exciting questions about how to use SAVs as feature embeddings for other tasks.
@@ Acknowledgements.
Here we provide additional information about additional experimental results, qualitative examples, implementation details, and datasets. Specifically, Section 7 provides more experiment results, Section 8 provides additional implementation details, and Section 9 provides qualitative visualizations to illustrate our approach.
7 Additional Experiment Results
We begin by presenting several additional ablations (Section 7.1) that further demonstrate the benefits of our SAVs approach. We also present additional results (Section 7.2) on BLINK Splits.
7.1 Additional Ablations
@@
7.2 Additional Results
(a) Impact of Token Position
MHB NB ES
Last 80.8 35.1 86.7
Middle 49.8 2.4 82.7
First 49.4 0 24.9
@@ 7.2 Additional Results
Table 5: SAV Additional Results. We perform several additional experiments to demonstrate different properties and capabilities of SAVs. In particular, we evaluate the effectiveness of our method (a) when selecting attention vectors from different tokens, (b) on language-only tasks, and (c) when using it in an online learning setting. Note: MHB represents MHaluBench, NB represents NaturalBench Group Score, ES represents EuroSAT, and O.L. represents online learning.
@@
Token position selection. Because the last-token of a sequence in a decoder-only LMM attends to all of the prior tokens in an input sequence, it is natural to extract SAVs from the heads of the last token. However, to validate this intuition, we compare the performance SAVs to extract sparse vectors from other tokens (first, middle, and last). Overall, our results in Table 5(a) show that the last token is the best option for selecting heads for SAVs.
@@
8 Additional Implementation Details